This is a differential equations course, please answer all the questions.

Find the general solution.

1 1 6. x(t) = 1 2 -8 -1 :(t) 1 -5 -3 3 2 4 7. z(t) = 2 0 2 *(t) 4 2 3 1 -1 2 8. :(t) 2 11. It) 3.-0 9. z(t) = (3 =) =(8)
Answer #1

1 - 6 x(+)- 2 211) 8 -5 -3 The eigen values of the co-efficient matrix is geven by i 1-3 2 1- -1 =0 -8 -5 -3-71 2 +4 -0 (1-124 = - 3, 2х, 4 z - я [zx he eigen veetor , у я 2Х, dat X=1 1. V - 2. fo1 2,2 -2 the eigen vector is given by, 4 1+2 — — 2for Az = 2 the eigen vector is given by, 21 1-2 - 1 2 1-2 -8 - 5 -3-2 23 -74 + x2 +33 = 0 274 -74 -X2=0 - 8x4 -58 -58z = 0 =>